OpenStudy (anonymous):

d + 3n = 1 solve for n

OpenStudy (ajprincess):

1.Subtract d both sides 2.Divide by 3 both sides.

OpenStudy (anonymous):

ajprincess: I think that is what I did and I got: n = 1 - d/3. But jiji501 has something different. Am I wrong?

OpenStudy (ajprincess):

yup. d + 3n = 1 1.Subtract d both sides d+3n-d=1-d 3n=1-d 2.Divide by 3 both sides. 3n/3=(1-d)/3 n=(1-d)/3 n=1/3-d/3 Nw is it clear?

OpenStudy (anonymous):

I did the exact same thing as you.... but you have your answer different than I do. I have n = 1-d/3 as one fraction and you have n = 1/3 - d/3 as two fractions I don't understand why you split it into 2 fractions.. why is that?

OpenStudy (ajprincess):